英語 (English)

基盤英語 (Basic English) (工 ((化 A)1 年))

早内-プリングルジュディス · 非常勤講師 1 単位前期金 5 ・ 6

(平成 19 年度以前の授業科目: 『基盤英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (1)』 )

【授業の目的】To give students opportunities to develop their English abilities through the exploration and discussion of a variety of themes. Students will develop their real-life communicative skills while expanding their vocabulary,grammar ability and fluency. A textbook will be used along with other media forms. Practice in the four skills reading, writing, listening and speaking will be provided during class and through homework.

【授業の概要】 Students will be expected to speak and write only in English during class. They will be expected to work not only in class, but also outside. There will be homework and they are expected to do it even if they are absent from class. The following class schedule is tentative and only shows the general contents of the textbook used in this course. It may be modified due to the size of the class or other factors.

【キーワード】 four skills， fluency 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】In today’s globalized society, English has become the common

language of communication and is necessary to gain more knowledge of world events, and process information. This class requires students to put their skills into practice, and in the process increase fluency and confidence in using English while also increasing vocabulary and polishing grammatical skills.

【授業の計画】 1. General Class Guidance; What is engineering?

2. Units 5 & 6: What is it made from?

3. Units 7 & 8: A picture is worth a thousand words; Getting away from it all

4. Units 9 & 10: The car’s a star; Made by hand

5. Presentation 1: Invention

6. Unit 11: Did you read the instructions?

7. Units 12 & 13: Safety first; Are you sitting comfortably?

8. Mid-term Evaluation; Mini Presentation 2

9. Units 14 & 15: Small is beautiful; Big is best 10. Unit 18: Gadgets

11. Units 19 & 20: Bridging the gap; Bridge disasters

12. Units 21 & 22: Ancient structures; Do you understand?

13. Units 23 & 24: How do you say? Where have I heard that name before?

14. Units 25 & 26: What’s your number?; It’s all just numbers; Presentation 3:A Fantastic Product

15. Presentation 3: A Fantastic Product: Part 1

16. Presentation 3: A Fantastic Product: Part 2; Final Exam

【教科書】Lindsay White, Engineering Workshop, Oxford University Press

ISBN0-19-438827-1 【参考書等】授業中に指示します．【成績評価の方法】 The Mid-term and Final evaluations are an important component of this course, but weekly attendance, active participation in class work, and homework completion are even more important. Tentative coursework evaluation: Weekly class participation, assignment, and homework (60%), Mid-term and Final Evaluation (40%).

【再試験の有無】 Allowable if the student has attended more than 2/3 of the classes

and has a score of 50% or higher. 【受講者へのメッセージ】 Make the most of all opportunities to expand your horizons, and

especially when communicating in another language.

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220851 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】

⇒ 早内-プリングル (juditheph@yahoo.com) ���� (オフィスアワー: By appointment)

English

Basic English (工 ((化 A)1 年))

Judith Hayauchi · Part-time Lecturer 1 unit 前期金 5 ・ 6

(平成 19 年度以前の授業科目: 『基盤英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (1)』 )

Target⟩ To give students opportunities to develop their English abilities through the exploration and discussion of a variety of themes. Students will develop their real-life communicative skills while expanding their vocabulary,grammar ability and fluency. A textbook will be used along with other media forms. Practice in the four skills reading, writing, listening and speaking will be provided during class and through homework.

Outline⟩ Students will be expected to speak and write only in English during class. They will be expected to work not only in class, but also outside. There will be homework and they are expected to do it even if they are absent from class. The following class schedule is tentative and only shows the general contents of the textbook used in this course. It may be modified due to the size of the class or other factors.

�������⟩ four skills, fluency Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ In today’s globalized society, English has become the common language of

communication and is necessary to gain more knowledge of world events, and process information. This class requires students to put their skills into practice, and in the process increase fluency and confidence in using English while also increasing vocabulary and polishing grammatical skills.

Schedule⟩

1. General Class Guidance; What is engineering?

2. Units 5 & 6: What is it made from?

3. Units 7 & 8: A picture is worth a thousand words; Getting away from it all

4. Units 9 & 10: The car’s a star; Made by hand

5. Presentation 1: Invention

6. Unit 11: Did you read the instructions?

7. Units 12 & 13: Safety first; Are you sitting comfortably?

8. Mid-term Evaluation; Mini Presentation 2

9. Units 14 & 15: Small is beautiful; Big is best 10. Unit 18: Gadgets

11. Units 19 & 20: Bridging the gap; Bridge disasters

12. Units 21 & 22: Ancient structures; Do you understand?

13. Units 23 & 24: How do you say? Where have I heard that name before?

14. Units 25 & 26: What’s your number?; It’s all just numbers; Presentation 3:A Fantastic Product

15. Presentation 3: A Fantastic Product: Part 1

16. Presentation 3: A Fantastic Product: Part 2; Final Exam

Textbook⟩ Lindsay White, Engineering Workshop, Oxford University Press

ISBN0-19-438827-1 Reference⟩ 授業中に指示します． Evaluation Criteria⟩ The Mid-term and Final evaluations are an important component of this course, but weekly attendance, active participation in class work, and homework completion are even more important. Tentative coursework evaluation:Weekly class participation, assignment, and homework (60%), Mid-term and Final Evaluation (40%).

Re-evaluation⟩ Allowable if the student has attended more than 2/3 of the classes

and has a score of 50% or higher. Message⟩ Make the most of all opportunities to expand your horizons, and

especially when communicating in another language.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220851 Contact (Office-Hour, Room, E-mail)⟩

⇒ Hayauchi (juditheph@yahoo.com) ���� (Office Hour: By appointment)